Tom is driving from Dallas to Houston, which is 362,000 m away. He drives with an average speed of 1010 m/s. How long does it take him to get to Houston?

Grade:11

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Time required = distance/ speed

= 362000/ 1010 = 358.4158 seconds

Regards
Vikas TU
4 years ago
Dear studentTime required = distance/ speed= 362000/ 1010 = 358.4158 seconds
=> 358.4158 / 60* 60 = 0.099 hour 
Which is approximately = 6 minutes 
Hope this helps